# Behavioral and Topological Heterogeneities in Network Versions of Schelling’s Segregation Model

### v3

@@
Department of Systems Science and Industrial Engineering Binghamton University
@@ 1   Abstract
Agent-based models of residential segregation have been of persistent interest to various research communities since their origin with James Sakoda [1] and popularization by Thomas Schelling [2]. Frequently, these models have sought to elucidate the extent to which the collective dynamics of individual preferences may cause segregation to emerge. This open question has sustained relevance in U.S. jurisprudence. Previous investigation of heterogeneity of behaviors (preferences) by Xie & Zhou [3] has shown reductions in segregation. Meanwhile, previous investigation of heterogeneity of social network topologies by Gandica, Gargiulo, and Carletti [4] has shown no significant impact to observed segregation levels. In the present study, we examined the effects of the concurrent presence of both behavioral and topological heterogeneities in network segregation models. Simulations were conducted using both Schelling’s and Xie & Zhou’s preference models on 2D lattices with varied levels of densification to create topological heterogeneities (i.e., clusters, hubs). Results show a richer variety of outcomes, including novel differences in resultant segregation levels and hub composition. Notably, with concurrent increased representations of heterogeneous preferences and heterogenous topologies, reduced levels of segregation emerge. Simultaneously, we observe a novel dynamic of segregation between tolerance levels as highly tolerant nodes take residence in dense areas and push intolerant nodes to sparse areas mimicking the urban-rural divide.
2   Acknowledgement
A Thesis has previously been published [5].
@@ 3   Introduction
Residential segregation is a persistent topic of discussion in the social sciences whose investigation is frequently enriched by computational models, particularly agent-based models (ABMs) with network structure [1, 2, 3, 4, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Residential segregation appears to be a robust and resilient phenomenon with many possible contributing factors at multiple scales [8]. This work focuses primarily on one of the two traditions of segregation theory identified by Fossett [9]: “individual preferences,” which asserts that segregation is an emergent property arising from the collective dynamics of individuals and their choices in mostly-free housing markets. Identifying and understanding factors and behaviors which contribute to residential segregation has substantial sociopolitical significance. In the decades following Brown v. Board of Education, a legal debate surrounding de jure vs. de facto segregation with proponents of the de facto perspective citing the role of such collective dynamics of individual preferences as the genesis of de facto segregation [18]. Given the long-standing view that government should restrict interventions to cases where de jure segregation is evident, it is clearly important to understand the extent to which de facto segregation might (or might not) emerge as the result of the collective dynamics of individual preferences. Recent works in complex systems and network sciences [19, 20, 21, 22] have highlighted the importance of heterogeneity in network models of social problems, showing clearly that certain emergent, system-level phenomena may only arise with sufficient heterogeneity of agent behaviors and/or characteristics. A review of the literature identified many investigations of component heterogeneity in network models of segregation. Notably, Xie and Zhou [3] and Gandica, Gargiulo, & Carletti [4] examined heterogeneous preferences and heterogeneous topology, respectively. The former obtained reductions in resultant segregation when heterogeneous preferences were represented while the latter showed no substantial reduction in segregation when heterogeneous topology was represented. No previous studies of combined dimensions of heterogeneity were identified. Recognizing the possibility that multiple dimensions of heterogeneity may be required to observe additional phenomena, this study attempts to examine the impact of the combination of behavioral and topological heterogeneities in network models of residential segregation. A baseline model is constructed following Xie & Zhou’s model [3]. The model is then elaborated in two ways: first to introduce heterogeneity of preferences and second to introduce heterogeneity of topology. The results of these implementations are illustrated, observables are characterized, and simulations are enumerated. Finally, results, novel behaviors, and opportunities for future work are discussed.
@@ 4.1   Sakoda and Schelling
In 1971, Sakoda [1] and then Schelling [2], proposed lattice-based models of social interaction and, more specifically, segregation. These models can be considered some of the earliest instances of agent-based models (ABMs). While Sakoda’s work faded into relative obscurity, Schelling’s gained great renown [23]. Importantly, these works demonstrated the emergence of macro-level (population-level) phenomena that resulted from micro-level (individual) interactions with neighbors. Both showed that segregation was one of the macro-level properties which could emerge. Schelling’s investigation of segregation was more extensive. It included a detailed discussion of the dynamics of his “spatial proximity” model. Like Sakoda, Schelling’s model randomly arranged individuals and vacancies on a 2D lattice. Each individual was assigned a tolerance threshold of 0.50, that is individuals are unhappy when the neighborhood proportion of unlike neighbors exceeds 50 percent. Then, without particular order, individuals are selected for transfer to a tolerable vacant location. This procedure is repeated until all individuals are satisfied, or until no additional viable moves are available. The results were striking. Even without overt segregationist preferences, the collective dynamics of individuals’ preferences to avoid minority status resulted in near-total segregation. Additionally, Schelling recognized that varying tolerances (preference schedules) and increasing neighborhood sizes could generate a greater variety of results, including reduced segregation, in his “bounded-neighborhood” model. This model was not, however, concerned with the properties of the configurations of individuals within the neighborhood. Rather, it only considered whether they chose to remain within or exit. In the years after Schelling, substantial replication, extension, and discussion of the model took place [7]. Despite Schelling’s acknowledgment of his model’s crudity, it gained importance in the 1980s and 1990s with debates concerning the genesis and perpetuation of segregation. On one side were proponents of the notion that segregation in the United States was primarily caused and perpetuated by discriminatory housing policies. The other side suggested that could be caused by the collective dynamics of ethnic preferences aided by economic disparities [7]. The suggestion that segregation exists due to such collective dynamics has even been an important factor in U.S. court decisions as legal remedies to racial injustices often require evidence of de jure causes [18]. As this controversy unfolded, several researchers examined and elaborated Schelling’s work.
@@ 4.3   Parameter Elaboration
Abella, San Miguel, & Ramasco [13] incorporated an aging effect. Their model modifies an agent’s probability of migration by making it inversely proportional to time spent in a satisfactory neighborhood. The authors examined several homogenous tolerance regimes. The addition of the aging effect led to the surprising result that segregation was maintained even at higher tolerance levels.
Gambetta, Mauro, & Pappalardo [15] introduced mobility constraints. In their model, preference for nearby or distant candidate vacancies was controlled by a distance exponent, $\beta$. For $\beta>0$, more distant locations are preffered. For $\beta<0$, nearby locations are preferred. In addition, they incorporated a ”relevance” effect, controlled by a relevance exponent, $\alpha$, where $\alpha>0$ indicates a preference for migration toward the center of the grid. These effects had significant impacts on behavior both independently and in concert. The distance model alone showed increased segregation for $\beta>0$ and decreased segregation for $\beta<0$. The relevance model alone showed modestly increased segregation for $\alpha>0$. Increases in $\alpha$ were moderated when $\beta<0$ as the preference for nearby locations counteracted the relevance effect.
@@
4.4   Game-Theoretic Approach
@@
Zhang [11] translated Schelling’s “spatial proximity” model to a spatial game-theoretic model. In this, Zhang presumed asymmetrical tolerances between groups and added a simple housing market, again employing a 2D lattice. As prices in the housing market responded to demand, the asymmetry of tolerances played an important role. Housing for the group with more exclusionary tolerances became scarcer. Additionally, the quantity of housing units that were unsuitable for the exclusionary group began to exceed demand. The result was a significant disparity in prices based on neighborhood composition. Zhang showed that in the absence of such market influences, asymmetry of tolerances alone could not explain the genesis of segregation on the lattice. In a separate paper, Zhang [12] employed his game-theoretic approach on a 2D lattice to illustrate an extreme example: the genesis of segregation when all agents prefer total integration. This time, no vacancies were permitted on the lattice. Again, the asymmetry of tolerances was important. For each group, Zhang established a utility function with maximum payoff when there was a perfect mix of neighbors. Secondarily, when a perfect mix of neighbors was not available, agents preferred to avoid minority status. Zhang showed that this asymmetry of tolerances makes perfect mixing an unstable attractor. Whenever an agent finds a trading partner, the payoff gained by the former will always be outweighed by the penalty to the latter.
4.5   Metapopulation Model
From 2015 to 2018, Gandica, Gargiulo, & Carletti [4, 10] elaborated Schelling’s spatial proximity model using a metapopulation framework. Initially in [4], they constructed regular 1- and 2-D lattices. Each node on the lattice could house a population of individuals with some limit L. Thus, for each individual living in the ith node, its neighborhood size, $s_{i}$, the level of population in i plus the sum of the levels, l, of population in each neighbor, j. $s_{i}=\sum_{j\in N_{i}}l_{j}$, where $j\in N_{i}$ signifies the set of nodes j in the neighborhood $N_{i}$. The authors conducted simulations using several homogeneous tolerance thresholds, $\epsilon$. For $\epsilon\leq 0.5$, at the node level, total segregation was observed. It is important to note that the proposed metapopulation model is analogous to the instantiation of a heterogeneous topology superimposed on a lattice structure. To clarify, let L, the node population limit be 4. Then, the set of possible node population levels, $P=\left\{0,1,2,3,4\right\}$. For a regular 2D lattice with von Neumann neighborhoods, the range is 1 to 20, or $[\max(1,(4*\min{P})),(5*\max{P})]$. This is equivalent to adding edges to a 1- or 2-D lattice populated by single occupant nodes to achieve the same heterogeneity of neighborhood sizes, except that the method proposed by Gandica, Gargiulo, & Carletti increases the number of individual neighborhoods.
@@ 4.6   Heterogeneous Topology
Later Gandica, Gargiulo, & Carletti [10], elaborated their earlier work on 2D lattices to examine the effect of varying network structures while employing the metapopulation model. While maintaining an average node-neighborhood size, k=4, small-world (Watts-Strogatz [25]), random (Erdos-Renyi [26]), and scale-free (Barabasi-Albert [27]) networks were generated. Given the same node-level population limit, L, a greater variety of neighborhood sizes is possible. Results from this inquiry showed no qualitative difference in asymptotic averaged node magnetization, $\langle u\rangle$, behavior, where
@@
Similar to [10], but without use of a metapopulation model, Fagiolo , Valente, & Vriend [14] considered six distinct network topologies: 2D boundary-less lattices with von Neumann and Moore neighborhoods, regular and random non-directed graphs, and scale-free and small-world networks. The authors examined the impact of these topologies on the segregation process. Homogenous tolerance thresholds were employed. Consistent with [10], the results showed that the segregation process was not significantly impacted by the network topology.
@@ 5.1   Baseline Model
The baseline network model topology is a 32 by 32 regular lattice with edges connecting von Neumann neighborhoods (see Figure 2) and a closed boundary condition. As in Xie & Zhou [3], 15% of nodes are reserved as excess housing. The remaining 85% of nodes are randomly assigned either a red or blue occupant. As in Schelling [2], each occupant is assigned an identical tolerance threshold, $\epsilon$. We set $\epsilon=0.3962$, indicating a tolerance for at most  39.62% opposite colored neighbors. To draw a more direct comparison, this $\epsilon$-value is obtained by taking the mean tolerance threshold from the populations of agents generated using tolerance schedules as described in the next section. Again, mirroring Xie & Zhou’s [3] baseline model, at time t, the current neighborhood proportion of dissimilar neighbors for the ith occupant in the $j^{th}$ neighborhood in $G$, $d_{ijt}$, is given as
$d_{ijt}=\begin{cases}\frac{n_{d_{i,t}}}{N_{jt}},\&\text{if }N_{jt}>0\\ 0,\&\text{if }N_{jt}=0\end{cases}$ (2)
where $n_{d_{i}t}$ is the number of dissimilar neighbors and $N_{jt}$ is the total number of neighbors in the $j^{th}$ neighborhood in $G$. At each time step $t$, $d_{ijt}$ is calculated for each individual $i$ in the set of occupants $O$ and the set of candidate occupants at time $t$, $C_{Ot}=\left\{i\mid i\in O,d_{ijt}\in\epsilon_{i}\right\}$, is constructed where $j$ is the individual’s current location. A candidate occupant $i\in C_{Ot}$ is randomly selected. For all vacant nodes at time $t$, $v\in V_{t}$, neighborhood composition is calculated to create a list of candidate vacancies, $C_{Vt}=\{j\mid j\in Vt,\ d_{ijt}\leq\epsilon_{i}\}$, from which a destination node, $j\in C_{Vt}$ is randomly selected. The candidate occupant at the $i^{th}$ node moves to the selected vacancy and leaves a vacancy in its place. If at any time $t$, $C_{Ot}={\emptyset}$, all nodes are satisfied, and no additional trades will be found. If at any time $t$, $C_{Vt}={\emptyset}$, the occupant at the selected node, $i\in C_{Ot}$, is unable to locate a satisfactory destination and remains in place.
5.2   Heterogeneous Tolerances
Following Xie & Zhou [3], agents are provided with heterogeneous tolerances (tolerance thresholds) aligned with the Guttman scale and rank-ordered logit model derived from Bruch & Mare’s [6] Detroit data. Tolerance thresholds were assigned by drawing values from a uniform distribution over a given interval: For 10.47% of individuals, $\epsilon_{i}$ fell within $[0.0,0.07)$; for 18.10% of individuals, $\epsilon_{i}$ fell within $[0.07,0.21)$; for 26.73% of individuals, $\epsilon_{i}$ fell within $[0.21,0.36)$; for 13.86% of individuals $\epsilon_{i}$ fell within $[0.36,0.57)$; for 26.59% of individuals, $\epsilon_{i}$ fell within $[0.057,1.00]$. For these individuals, the simulation procedure described in the previous section was implemented with $\epsilon_{i}$ replacing $\epsilon$. Figure 1 shows an example distribution of tolerance thresholds drawn from the constructed distribution. For the 4.25% of individuals in the Detroit data found not to conform to the Guttman scale, no static tolerance threshold was set. Instead, Xie & Zhou’s rank-ordered logit model (Eq. 4 in [3]) was implemented to determine the probability of transition to each candidate neighborhood. The transition destination, $j\in C_{Vt}$, is then randomly selected with probability, $\widehat{p_{jt}}$, given by the model:
[…]
The adoption of Xie & Zhou’s tolerance model increases the Shannon entropy of tolerance thresholds, $H_{\epsilon}$, from 00 in the homogenous Schelling case to at least $H_{\epsilon}\approx 5.102$. This approximation was calculated by discretizing assigned tolerance thresholds into 37 identically sized bins and allowing non-Guttman scale agents to occupy an additional distinct bin.

### v4

@@
School of Systems Science and Industrial Engineering Binghamton University
@@ 1   Abstract
Agent-based models of residential segregation have been of persistent interest to various research communities since their origin with James Sakoda and popularization by Thomas Schelling. Frequently, these models have sought to elucidate the extent to which the collective dynamics of individual preferences may cause segregation to emerge. This open question has sustained relevance in U.S. jurisprudence. Previous investigation of heterogeneity of behaviors (preferences) has shown reductions in segregation. Meanwhile, previous investigation of heterogeneity of social network topologies has shown no significant impact to observed segregation levels. In the present study, we examined the effects of the concurrent presence of both behavioral and topological heterogeneities in network segregation models. Simulations were conducted using both homogeneous and heterogeneous preference models on 2D lattices with varied levels of densification to create topological heterogeneities (i.e., clusters, hubs). Results show a richer variety of outcomes, including novel differences in resultant segregation levels and hub composition. Notably, with concurrent increased representations of heterogeneous preferences and heterogeneous topologies, reduced levels of segregation emerge. Simultaneously, we observe a novel dynamic of segregation between tolerance levels as highly tolerant nodes take residence in dense areas and push intolerant nodes to sparse areas mimicking the urban-rural divide.
@@ 2   Introduction
Residential segregation is a persistent topic of discussion in the social sciences whose investigation is frequently enriched by computational models, particularly agent-based models (ABMs) with network structure [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Residential segregation appears to be a robust and resilient phenomenon with many possible contributing factors at multiple scales [7]. Several authors have demonstrated significant interplay between racial segregation and wealth and income inequality [4, 5, 17, 18]. Others have developed supporting theoretical and simulation models to understand this relationship [19, 20, 21, 22, 23, 24, 25]. Identifying and understanding factors and behaviors which contribute to residential segregation has substantial sociopolitical significance. A legal debate over the causes of segregation emerged with the Brown v. Board of Education decision in 1954. Whether segregation is a de jure or de facto phenomenon remains unresolved. The government’s long-standing view is that it should restrict its interventions to cases where de jure segregation is evident.
Of the two traditions of segregation theory identified by Fossett [8]: “social distance” and “individual preferences,” this work focuses on the latter which asserts that segregation is an emergent property arising from the collective dynamics of individuals and their choices in mostly-free housing markets. Proponents of the de facto perspective cite the collective dynamics of individual preferences as the genesis of de facto segregation [26]. Given the current state of related jurisprudence, it is clearly important to understand the extent to which de facto segregation might (or might not) emerge as the result of the collective dynamics of individual preferences.
@@
Recent works in complex systems and network sciences [27, 28, 29, 30] have highlighted the importance of heterogeneity in network models of social problems, showing clearly that certain emergent system-level phenomena may only arise with sufficient heterogeneity of agent behaviors and/or characteristics. A review of the literature identified many investigations of component heterogeneity in network models of segregation. Notably, Xie and Zhou [3] and Gandica, Gargiulo, & Carletti [4] examined heterogeneous preferences and heterogeneous topology, respectively. The former obtained reductions in resultant segregation when heterogeneous preferences were represented while the latter showed no substantial reduction in segregation when heterogeneous topology was represented. No previous studies of combined dimensions of heterogeneity were identified. Recognizing the possibility that multiple dimensions of heterogeneity may be required to observe additional phenomena, this study attempts to examine the impact of the combination of behavioral and topological heterogeneities in network models of residential segregation. A baseline model is constructed following Xie & Zhou’s model [3]. The model is then elaborated in two ways: first to introduce heterogeneity of preferences and second to introduce heterogeneity of topology. The results of these implementations are illustrated, observables are characterized, and simulations are enumerated. Finally, results, novel behaviors, and opportunities for future work are discussed.
@@ 3.1   Sakoda and Schelling
In 1971, Sakoda [1] and then Schelling [2], proposed lattice-based models of social interaction and, more specifically, segregation. These models can be considered some of the earliest instances of agent-based models (ABMs). While Sakoda’s work faded into relative obscurity, Schelling’s gained great renown [31]. Importantly, these works demonstrated the emergence of macro-level (population-level) phenomena that resulted from micro-level (individual) interactions with neighbors. Both showed that segregation was one of the macro-level properties which could emerge. Schelling’s investigation of segregation was more extensive. It included a detailed discussion of the dynamics of his “spatial proximity” model. Like Sakoda, Schelling’s model randomly arranged individuals and vacancies on a 2D lattice. Each individual was assigned a tolerance threshold of 0.50, that is individuals are unhappy when the neighborhood proportion of unlike neighbors exceeds 50 percent. Then, without particular order, individuals are selected for transfer to a tolerable vacant location. This procedure is repeated until all individuals are satisfied, or until no additional viable moves are available. The results were striking. Even without overt segregationist preferences, the collective dynamics of individuals’ preferences to avoid minority status resulted in near-total segregation. Additionally, Schelling recognized that varying tolerances (preference schedules) and increasing neighborhood sizes could generate a greater variety of results, including reduced segregation, in his “bounded-neighborhood” model. In bounded-neighborhood models agents consider their membership in a discrete, finite neighborhood. This model was not, however, concerned with the properties of the configurations of individuals within the neighborhood. Rather, it only considered whether they chose to remain within or exit. In the years after Schelling, substantial replication, extension, and discussion of the model took place [6]. Despite Schelling’s acknowledgement of his model’s crudity, it gained importance in the 1980s and 1990s with debates concerning the genesis and perpetuation of segregation. On one side were proponents of the notion that segregation in the United States was primarily caused and perpetuated by discriminatory housing policies. The other side suggested that could be caused by the collective dynamics of ethnic preferences aided by economic disparities [6]. The suggestion that segregation exists due to such collective dynamics has even been an important factor in U.S. court decisions as legal remedies to racial injustices often require evidence of de jure causes [26]. As this controversy unfolded, several researchers examined and elaborated Schelling’s work.
@@ 3.3   Parameter Elaboration
Abella, San Miguel, & Ramasco [12] incorporated an aging effect. Their model modifies an agent’s probability of migration by making it inversely proportional to time spent in a satisfactory neighborhood. The authors examined several homogeneous tolerance regimes. The addition of the aging effect led to the surprising result that segregation was maintained even at higher tolerance levels.
Gambetta, Mauro, & Pappalardo [14] introduced mobility constraints. In their model, preference for nearby or distant candidate vacancies was controlled by a distance exponent, $\beta$. For $\beta>0$, more distant locations are preferred. For $\beta<0$, nearby locations are preferred. In addition, they incorporated a “relevance” effect, controlled by a relevance exponent, $\alpha$, where $\alpha>0$ indicates a preference for migration toward the center of the grid. These effects had significant impacts on behavior both independently and in concert. The distance model alone showed increased segregation for $\beta>0$ and decreased segregation for $\beta<0$. The relevance model alone showed modestly increased segregation for $\alpha>0$. Increases in $\alpha$ were moderated when $\beta<0$ as the preference for nearby locations counteracted the relevance effect.
@@
Several other computational modeling efforts have examined the interplay between segregation and wealth inequality. Benard & Willer showed that individual resource differences can enable segregation by status on a toroidal cellular automata model [19]. Sahasranaman & Jensen, in two papers, developed and extended a model of wealth-based segregation which well-approximated empirical evidence when individuals had a preference for neighborhoods in which their wealth was comparable to or greater than their neighbors [21, 22]. Gauvin et al. provide contradictory evidence [20]. That is, their model which also provided a reasonable approximation of empirical data showed that wealth-based segregation could occur when individuals prefer to live in neighborhoods in which their wealth was less than or comparable to their neighbors. Harting & Radi, used a model with heterogeneous tolerances and ethnic income disparities to show that when individuals make housing decisions based on economic factors in addition to ethnic preferences, income inequality may be a necessary condition for residential integration [23]. Bonakdar & Roos used heterogeneous incomes and incorporated an auction process to determine housing prices based on demand [24]. Their simulation results supported the idea that when individuals have a preference for similar neighbors, population sorting motivates housing prices. Vicario implemented a model in which agents are characterized by a dynamic wealth variable showing that, absent perturbations at the individual level, wealth segregation emerges [25].
3.4   Game-Theoretic Approaches
@@
Zhang [10] translated Schelling’s “spatial proximity” model to a spatial game-theoretic model. In this, Zhang presumed asymmetrical tolerances between groups and added a simple housing market, again employing a 2D lattice. As prices in the housing market responded to demand, the asymmetry of tolerances played an important role. Housing for the group with more exclusionary tolerances became scarcer. Additionally, the quantity of housing units that were unsuitable for the exclusionary group began to exceed demand. The result was a significant disparity in prices based on neighborhood composition. Zhang showed that in the absence of such market influences, asymmetry of tolerances alone could not explain the genesis of segregation on the lattice. In a separate paper, Zhang [11] employed his game-theoretic approach on a 2D lattice to illustrate an extreme example: the genesis of segregation when all agents prefer total integration. This time, no vacancies were permitted on the lattice. Again, the asymmetry of tolerances was important. For each group, Zhang established a utility function with maximum payoff when there was a perfect mix of neighbors. Secondarily, when a perfect mix of neighbors was not available, agents preferred to avoid minority status. Zhang showed that this asymmetry of tolerances makes perfect mixing an unstable attractor. Whenever an agent finds a trading partner, the payoff gained by the former will always be outweighed by the penalty to the latter. More recently, Grauwin et al. [33] built upon Zhang’s work by employing evolutionary game theory to develop a potential function for analysis of Schelling’s bounded-neighborhood model. Using this potential function, they analyzed stationary configurations generated by three different utility functions: linear utility, Schelling’s original function, and asymmetric peaked utility functions. They showed that symmetric functions do not generate segregation, but, as in Zhang [11], any preference for majority status, even when individuals primarily prefer mixed environments, generates segregation. Their analytical tool enables exploration of an array of utility functions, including those rooted in empirical data.
@@ 3.5   Heterogeneous Topology
From 2015 to 2018, Gandica, Gargiulo, & Carletti [4, 9] elaborated Schelling’s spatial proximity model using a metapopulation framework. Initially in [4], they constructed regular 1- and 2-D lattices. Each node on the lattice could house a population of individuals with some limit $L$. Thus, for each individual living in the $i^{th}$ node, its neighborhood size is the level of population in $i$ plus the sum of the levels, $l$, of population in each neighboring node. The authors conducted simulations using several homogeneous tolerance thresholds, $\epsilon$. For $\epsilon\leq 0.5$, at the node level, total segregation was observed. It is important to note that the proposed metapopulation model is analogous to the instantiation of a heterogeneous topology superimposed on a lattice structure. This is equivalent to adding edges to a 1- or 2-D lattice populated by single occupant nodes to achieve the same heterogeneity of neighborhood sizes, except that the method proposed by Gandica, Gargiulo, & Carletti increases the number of individual neighborhoods.
Later Gandica, Gargiulo, & Carletti [9], elaborated their earlier work on 2D lattices to examine the effect of varying network structures while employing the metapopulation model. While maintaining an average node-neighborhood size, k=4, small-world (Watts-Strogatz [34]), random (Erdos-Renyi [35]), and scale-free (Barabasi-Albert [36]) networks were generated. Given the same node-level population limit, L, a greater variety of neighborhood sizes is possible. Results from this inquiry showed no qualitative difference in the chosen metric for long-term average node tendency, asymptotic averaged node magnetization, $\langle u\rangle$, in Eq. 1.
@@
where $n_{i}^{A}$ and $n_{i}^{B}$ are the number of A and B individuals in the $i^{th}$ node, respectively, and $N$ is the total number of nodes.
Similarly, but without use of a metapopulation model, Fagiolo, Valente, & Vriend [13] considered six distinct network topologies: 2D boundary-less lattices with von Neumann and Moore neighborhoods, regular and random non-directed graphs, and scale-free and small-world networks. The authors examined the impact of these topologies on the segregation process. Homogeneous tolerance thresholds were employed. Consistent with [9], the results showed that the segregation process was not significantly impacted by the network topology. In 2023, Su & Zhang [37] implemented Schelling’s metapopulation model on a star-shaped network of $Q$ blocks. Each block with a population limit of $M$ agents was connected to a central hub. In addition, individual’s migration was constrained to adjacent blocks. This construction substantially suppressed the segregation phase.
@@ 4.1   Baseline Model
The baseline network model topology is a 32 by 32 regular lattice with edges connecting von Neumann neighborhoods (see Figure 1) and a closed boundary condition. As in Xie & Zhou [3], approximately 15% of nodes are reserved as excess housing. The remaining nodes are randomly assigned either a red or blue occupant. As in Schelling [2], each occupant is assigned an identical tolerance threshold, $\epsilon$. We set $\epsilon=0.3962$, indicating a tolerance for at most  39.62% opposite colored neighbors. To draw a more direct comparison, this $\epsilon$-value is obtained by taking the mean tolerance threshold from the populations of agents generated using tolerance schedules as described in the next section. Again, mirroring Xie & Zhou’s [3] baseline model, at time t, the current neighborhood proportion of dissimilar neighbors for the $i^{th}$ occupant in the $j^{th}$ neighborhood in $G$, $d_{ijt}$, is given as
$d_{ijt}=\begin{cases}\frac{n_{d_{i,t}}}{N_{jt}},&\text{if }N_{jt}>0\\ 0,&\text{if }N_{jt}=0\end{cases}$ (2)
where $n_{d_{i}t}$ is the number of dissimilar neighbors and $N_{jt}$ is the total number of neighbors in the $j^{th}$ neighborhood in $G$.
At each time step $t$, $d_{ijt}$ is calculated for each individual $i$ in the set of occupants $O$ and the set of candidate occupants at time $t$, $C_{Ot}=\left\{i\mid i\in O,d_{ijt}>\epsilon_{i}\right\}$, is constructed where $j$ is the individual’s current location and $\epsilon_{i}$ is its tolerance threshold. A candidate occupant $i\in C_{Ot}$ is randomly selected. For all vacant nodes at time $t$, $v\in V_{t}$, neighborhood composition is calculated to create a list of candidate vacancies, $C_{Vt}=\{j\mid j\in Vt,\ d_{ijt}\leq\epsilon_{i}\}$, from which a destination node, $j\in C_{Vt}$ is randomly selected. The candidate occupant at the $i^{th}$ node moves to the selected vacancy and leaves a vacancy in its place. If at any time $t$, $C_{Ot}={\emptyset}$, all nodes are satisfied, and no additional trades will be found. If at any time $t$, $C_{Vt}={\emptyset}$, the occupant at the selected node, $i\in C_{Ot}$, is unable to locate a satisfactory destination and remains in place.